Two point charges are brought together as shown below. Calculate the electric potential energy between the two charges if q1=2 nC and q2= 3 nC.


1
Expert's answer
2022-03-22T15:47:57-0400

Answer

the electric potential energy between the two charges is given by

"U=\frac{kq_1q_2}{r}\\=\frac{9*10^9*2*3*10^{-18}}{r}\\=\frac{54*10^{-9}}{r}"

Putting the distance find the desired results